OpenStudy (anonymous):

2x - 1 + 3(x - 5) = -2(2x + 1) Please give an Explanation and Answer

OpenStudy (anonymous):

2x-1+3(x-5)=-2(2x+1) multiply out 2x-1+3x-15=-4x-2 Combine like terms 5x-16=-4x-2 9x=14 x=14/9

OpenStudy (anonymous):

Thank you! Have a great day!